Tech Staff Meeting Notes June 3, 2009

The top priority is always repair of down stations.

The task list updated February 6, 2009 needs to be updated again. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


I40 Weigh Station Install
AR Network survey
Ft Payne station
CVVA trigger upgrades
SM sample rate fix

News & Highlights

The Fiscal Fish Fry will be June 26. Mac is in charge. Jenny is collecting the "donations" (Michelle Smith is collecting them while Jenny is on vacation).

Mitch and Chris spent quality time with the auditor. Its an internal audit and should be viewed as an opportunity to improve things. We want to find out what procedures we're not doing or need to fix before there's a problem and a review by an objective audit is the best way to do that. Its internal so any flaws stay internal to the University and gives us time to address them. Please cooperate if you're asked any questions.

We need to do a better job tracking our inventory and not just for response purposes. We need to better accomodate the annual inventory needs of both the UoM and USGS for hardware in our custody. There will be an Inventory Task Force (the ITF) charged with developing and maintaining a system that meets the needs (and policies) of the UoM and the USGS. It shall include model nubmers, property tag numbers (both UoM and USGS), serial numbers, location, custodian, etc. The task for will meet quarterly and conduct the inventory at least annually. Members of the task force are Jenny Spikes, Al Sedaghat, Jim Bollwerk, and Deshone Marshall.


Short Period & Broadbands

Holly's list:


Muted Amplitude

Isis Oxidation problem


Gain Channel problem

Broadband down


PWLA There is still significant telemetry interference that seems to be greatest during the work day. Chris replaced filter hardware and it looks much improved--or we're just having some very nice (and low humidity) weather with no inversions.

Several of the stations can get improved telemetry by transmitting directly to the node. The original configuration changed enough so that there is room in the band. Jim will investigate.

NODE Tasks

MKTA: Keep getting tflr messages. It could be sky view causing frequent single satellite lock. Al will adjust env_alarm.d to failure duration to 240 seconds. Jim will check the steiner alarm gizmo.

LNXT: Still need to tie generator status alarms into lnxtworm

NMAD: Repair barbed wire fence damaged by falling ice

AZTN: Upgrade PC and replace XLAK clock

Steve is still working with telecom to order dsl for the three NMSZ nodes and the new Boswell node. There are a couple snags. They said it isn't available at lnxt though we did have ISDN there at one time (the are similar protocols though ISDN can go a little farther than dsl).

Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected).


CVTN optical isolation. Need to upgrade to year 3 hardware.

RDST needs upgrade to year 3 hardware with router.

NAIT: Couldn't get the NWS cable modem to cooperate at NAIT (firewall?). Dean was going to ask Comcast about it and maybe get a new modem. Steve will follow up.

Need to complete firmware upgrade on CVVA.

Now that CMG5 triggers actually contain data, we need to tune the trigger level. Mitch sent explanatory info and guidelines to Al and Steve to tune.

SLU had trouble getting triggers from the 10sps datastream. They are having success creating 200sps triggers off the 100sps channel without excessive triggers (though I wonder what their levels are). The two newer CMG5TD's we got from ASL also don't trigger well off the 10sps datastream. Jim will investigate the feasibility of switching to the 100sps tap as the trigger source for the 200sps upgrade. Then Al will retune as above.

Most of the strong motion stations can run at 100 sps without any modifications. We will install a router on the few that have problems (RDST will be the test case).

Remove and repair recorders 1 and 6, Repair or remove and repair recorder 2. The recorder disks should be reformatted as a routine maintenance task next time access is available. The constant read/write to the fat file systems lead to fragmentation. Don't ya just love dos? Wireless link to # 5 is unstable. Wireless link to #7 has been down for several weeks, not sure of recorder status. Verify if recorders will work with windowsXP while we have several here at CERI.
We will have lane closures and access to a snooper truck in a couple weeks. Steve thinks its just the flash disks that need to be reformatted or replaced.

Installation is set for week of June 8. Still need dsl at Bosewell to get data flowing.

Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data.



RCGA charging circuitry at is old and primitive and in need of an upgrade.

CCRT - Needs a rebuild after a bad lightning strike.

GMG - Need battery box solar reg mounting plate and need to fuse the battery.


GFM - Fixed the bad channel but still needs some work to repair storm damage. The NC state park service is purchasing GFM. We already have a couple permissions from them and John initiated the paperwork for the GFM permission.

TRYN: the forest service wants to remove the tower. There will be a new 180' tower and we have permission to be on it. We also have permission to be in the building (rack mount only). Can we convert this station to AC?

SWET: This will turn into a "standard" CERI broadband. RT130 with trillium and episensor. We can keep it wireless and solar.

CPCT: The aluminum enclosure with the GPS breakout board at CPCT is corroded, fragile, and needs replaced. Need to re-work lighting and pest chaser. John has a pest chaser and will use some landscape or pool hardware for the lighting. A new breakout and transient suppression enclosure is ready. Needs a locking tower box. This will also be upgraded to the standard bb.

FPAL John found a good host for a Ft Payne broadband. There is AC (or rather the host is willing to run power). The ISP is also willing to provide free bandwidth though this may not last long. There is a second possibility where we could telemeter to Jacksonville State (

CMG6 was installed and returned at FPAL. It was installed on a glass plate concreted to rock believed to be bedrock. The sensor was in a garbage bag and backfilled with sand. The site is at the edge of woods with a large tree about 10 feet away. There was also a nearby tower though we didn't see excessive noise from that nor did there appear to be pump noise in the data. It was sensor number 6240. Mitch noted that:

It looks good just taking a quick pql view of the z traces. It will take time for me to massage them into day long mseed files for processing with PDF. Were John and Chris there on 5/25? There're alot of transient and gaps that day. There's a nice blast on 5/27 at about 19:30. The Guatemala earthquake shows up nicely on 5/28 at about 08:28. There's a bunch of garbage, both spikes and longer period glitches, on the east channel that looks instrumental or vault related or something not real. The north channel looks good (a little windy 5/24-26).

Jim installed the FPAL CMG3 in the test vault and the data look good. PDF's are at


James has a very nice SOH website at

Comm to McCarty is out. It appears to be working at both ends but they still don't connect. James is waiting to hear back from the ISP. Steve noted that there once was a problem at the UM gateway for one of the nodes so that's also a possibility if nothing else is found.

MAIR and Troy need filters, etc cleaned (can hit Troy when replacing the GLAT sensor).

Need to contact new owners (the water district) of the building at Miss. Co. Airport to refresh the permission.

The next Davis brain dump needs to be scheduled.

AR Network

Did a recon at several sites and installed a CMG6 at Powhatan. The temp station is online as ASN3.HH?.AG.

Greg determined that current draw should be about 675mA at 12.0 VDC assuming a 70% transmit ratio for the freewave. Two 55-60 watt panels should suffice and to allow for temps down to 10 degrees F, we should use four 105ah batteries to get 14 days backup power.



Retrieve cabinet and cable from the Moscow extension facility.

What to do with the Steiner brain dump DVD's?

Mitch will put John's database in seisadm on smeagol and locally at CERI we can connect via gaia. Al with get with Deshone to find out how we can use vpn or something to connect remotely (needs this for both John's access, and for field crews).

Steve and Al are setting up a feed of Oklahoma data to be forwarded to the NEIC. The scream2scream connection didn't seem to be all that stable and OKGS could use a better way to archive events before they leave the buffer. Steve will help them get ew going and then we'll use the ew export. We need to make sure to add relevant channels to the NEIC export when available and contact Dan McManus dmcmanus) at USGS Golden when they're going.

Rotational sensor is feeding an earthworm but still has IRIG crosstalk problem to fix. Its fine on an Oscope hooked to the output of the mux and looks fine in LabView. Its bad when viewed with wave_viewer.

Apparently the person on campus retired, who handles selling vehicles. Mac is looking into how to do this ourselves. And he needs to QUIT STALLING AND GET THIS DONE PRONTO!! (we look like a used car lot).

Pink Palace is working with Gary and Michelle on a new perc-like display area. We'll pull the receiver etc, when they start pulling out the old display (or sometime earlier if an opportunity arises).

(Old note from January) Mitch briefly reviewed the desired 106x6 network of desirable TA stations to adopt. These are the 106 TA stations within the ANSS RSN footprint for the NSMZ and WVSZ and the ETSZ and CSZ. The "by 6" implies three components for broadband and three of strongmotion. Considerable funds must be secured for the capitalization costs and for the long term operation and maintenance.


3890 Power Chris is going to make a diagram for 3890. Good luck son.

3876 Power Chris is going to make a diagram for 3876. Good luck son.